Circular No. 6475

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1996 S1 (HOLT-OLMSTEAD) J. V. Scotti, University of Arizona, reports his recovery of comet

P/1990 R2 (= 1990k = 1990 XVIII) with the Spacewatch telescope at Kitt Peak. The comet appears essentially stellar with a slight hint of diffuseness, and is close to the prediction on MPC 23482 (ephemeris on MPC 26588).

1996 UT  α_{2000}  δ_{2000}  m_{1}
Sept.19.20204  21^{h}39^{m}56.^{s}19  −29^{o}50′18′′.2  20.2
19.21964  21 39 55.54  −29 50 10.1  20.1
19.23678  21 39 54.95  −29 50 02.9  20.1
20.20399  21 39 23.78  −29 42 54.9  20.6
20.22183  21 39 23.17  −29 42 48.2  20.5
20.23900  21 39 22.61  −29 42 39.2  20.4

NGC 6624 S. F. Anderson, B. Margon, and E. W. Deutsch, University of Wash-

ington; R. A. Downes, Space Telescope Science Institute; and R. G. Allen, University of Arizona, report: “A Hubble Space Telescope observation of the ultraviolet-bright counterpart to the 11-min binary x-ray source in the core of the globular cluster NGC 6624 (King et al. 1993, Ap.J. 413, L117) shows a highly-statistically significant ultraviolet flux modulation with an 11.5-min (± 0.2 min) period, consistent with the 685-s period of the known x-ray modulation. The ultraviolet amplitude is very large compared with the observed x-ray oscillations: x-ray variations are generally reported as 2–3 percent peak-to-peak, whereas our data show an amplitude of order 17 ± 4 percent in the range 125–250 nm. These Faint Object Spectrograph data also have spectral resolution of 0.7 nm, so later analysis may reveal spectral-dependent fluctuations in this amplitude. Arons and King (1993, Ap.J. 413, L121) have predicted periodic ultraviolet fluctuations in this shortest-known-period binary system, due to the cyclically changing aspect of the x-ray-heated face of the secondary star. However, this predicted modulation has never before been observed. Employing their formalism, which invokes a number of different physical assumptions, we infer a system orbital inclination between ∼ 35^{o} and 55^{o}. Additional observations of this exotic system are badly needed.”

1996 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green